OpenStudy (anonymous):

In the coordinate plane, the point A (−4, -1) is translated to the point A '(−5, 0) . Under the same translation, the points B( −1, 2) and C (0, 3_ are translated to B ' and C ' , respectively. What are the coordinates of B 'and C '?

OpenStudy (anonymous):

21.

OpenStudy (michele_laino):

if you call \[x \prime , y \prime\] the coordinates of the A' point, and with \[x,y\] the onesof the A point, you can find these equations: \[x \prime =x-1, y \prime = y+1\], now you have to substitute the coordinates of your points B and C

OpenStudy (anonymous):

I dont really get it !

OpenStudy (michele_laino):

for point B, for example ou have \[x=-1, y=2\], then, please try to calculate \[x \prime , y \prime \]

OpenStudy (michele_laino):

Please try using the above formulae which I wrote